OpenStudy (anonymous):

Curve C: r(t) = (cos(t), sin(t), t-cos(t)) Tangent line C through P = r(pi/2)?

OpenStudy (anonymous):

r'(t) = (-sin(t) , cos(t) , 1+sin(t) )

OpenStudy (anonymous):

Now you just need a position vector r(pi/2) = (0,1,pi/2)

OpenStudy (anonymous):

1+sin(pi/2) = 2 right? but now you have the equation r(t) = (0,1,pi/2) + t(-1,0,0) and then?

OpenStudy (anonymous):

but what is the the step after you find r(t) ...

OpenStudy (anonymous):

r'(pi/2) = (-1,0,2)

OpenStudy (anonymous):

\[\gamma (t) = (0,1, \frac{\pi}{2}) + t (-1,0,2) \]\[t \in \mathbb{R}\]

OpenStudy (anonymous):

called it gamma, because technically the curve was called r , so yeah.

OpenStudy (anonymous):

oke that make sense.. but than you will have 3 tangent lines? one for the x, y en z?

OpenStudy (anonymous):

The equation a line is always \[x(t) = a + bt\] Where "x" is a component vector, "a" is the position vector and "b" is the direction vector, which you "span" along. As it is called in linear algebra. t is the free parameter, which you vary through the real numbers.

OpenStudy (anonymous):

No, you have ONE line, in 3 dimensional space.

OpenStudy (anonymous):

Picture a supporting cable on a tower (assuming it is not loose). That is a line in three dimensional space.

OpenStudy (anonymous):

and similarly it works for as many dimensions as you like. The general equation of a line will always be the same.

OpenStudy (anonymous):

The only difference will be the number of components in the position and direction vectors.

OpenStudy (anonymous):

could you type in the final answer for me? (a) so i can understand it :)

OpenStudy (amistre64):

teh vector functions derives componentwise into a tangent vector

OpenStudy (amistre64):

if you want to know the line of the tangent, you simply attach the tangent vector to a point

OpenStudy (amistre64):

elec did a good derivative for you; not you just need to know its value at pi/2

OpenStudy (amistre64):

r'(t) = (-sin(t) , cos(t) , 1+sin(t) ) r'(pi/2): Tx = -sin(pi/2 Ty = cos(pi/2) Tz = 1+sin(pi/2)

OpenStudy (amistre64):

what is that then, <0,1,1> ?

OpenStudy (amistre64):

so we nee to attach that to the point at the vector functions values of pi/2

OpenStudy (amistre64):

i got my tangent points wrong .. confused my values

OpenStudy (anonymous):

where did you found <0,1,1> ?

OpenStudy (anonymous):

Tx=0 Ty=0 Tz=2 right?

OpenStudy (anonymous):

oops Tx = 1

OpenStudy (amistre64):

.... ok, the only way I see for you to discover this is for you to get your hands dirty :) so lets start from the beginning and see where it goes awry at How do we derive r(t) = (cos(t), sin(t), t-cos(t))

OpenStudy (amistre64):

my calculating of sin and cos was wrong because I confused the 2

OpenStudy (anonymous):

oke.. de derivative of r(t) is: r'(t) = ( -sin(t), cos(t), 1+ sin(t) )

OpenStudy (amistre64):

those are good, now what do they represent to us?

OpenStudy (anonymous):

its the i dont know the word in english >.< but its the slope?

OpenStudy (amistre64):

slope is fine, but slope for a vector is just its component parts: r'(t) = Tangent vector <x(t),y(t),z(t)>

OpenStudy (amistre64):

so lets calculate correctly the component parts of our tangent vector

OpenStudy (anonymous):

owh jah ofcource the tangent vector =.=

OpenStudy (anonymous):

oke i have to fill in the pi/2 for t right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

oke lets do that

OpenStudy (anonymous):

r'(pi/2) = ( -sin(pi/2), cos(pi/2), 1+ sin(pi/2) ) r'(pi/2) = ( -1, 0, 2 )

OpenStudy (amistre64):

I get, hopefully :) Tx = -sin(pi/2 = -1 Ty = cos(pi/2) = 0 Tz = 1+sin(pi/2) = 1+1 = 2 T<-1,0,2> yay we match!

OpenStudy (anonymous):

hahah score ^^

OpenStudy (amistre64):

now that we have a vector all we need is a point to attach it to and a scalar to strectch it to infinity by and we get a line

OpenStudy (amistre64):

out point is just the original vector functions at pi/2

OpenStudy (amistre64):

*our point

OpenStudy (anonymous):

oke so: r(t) = (0,1,pi/2)

OpenStudy (amistre64):

thats a good point; now lets attach our vector to it :) ill use "s" for a scalar x = 0 + s*Tx y = 1 + s*Ty z = pi/2 + s*Tz

OpenStudy (amistre64):

T<-1,0,2> x = -s y = 1 + 0s z = pi/2 + 2s x = -s y = 1 z = pi/2 + 2s is simplified if i see it right

OpenStudy (anonymous):

oke... and what is the next step.. or is this the answer?

OpenStudy (amistre64):

this is what we needed, we defined a line by a point and any scaled version of its vector

OpenStudy (amistre64):

there are 2 forms that I know of, this one is the parametric; the other is similar and seldom used tho

OpenStudy (anonymous):

ahhh oke... the assignment said find the tangent line. so i thought there i just 1 line... stupid me

OpenStudy (amistre64):

:) oh its one "line" its just defined by its seperate components

OpenStudy (amistre64):

there is really no good way to define a line in 3 coords other than this

OpenStudy (anonymous):

ahhh oke :) thank you honey <3 i have another question could you watch if i do it right?

OpenStudy (amistre64):

if need be, since tis line acts nicely in one plane; we could define it like normal, we would just have to amend the equation by saying y = 1 so that it resides in the y=1 plane

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

r(t) = (1 - t^2, 2t) point p= (0,2)

OpenStudy (anonymous):

i also have to find the tangent line, but the t is not defined

OpenStudy (anonymous):

i know that r'(t) = ( -2t, 2 )

OpenStudy (amistre64):

the t is defined, you just have to unbury it from the given point

OpenStudy (anonymous):

(0,2) + t(-2t,2) is this step correct?

OpenStudy (amistre64):

not quite, :)

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

youve got it correct up to r'(t), but the tangent vector needs to be defined at (0,2) so we have to discover the secret to "t"

OpenStudy (amistre64):

tell me what x equals in the given r(t)

OpenStudy (anonymous):

x = -2t y = 2 point x = 0 en y = 2 so x = -2t 0 = -2t t=0?

OpenStudy (anonymous):

oops i took the derivative

OpenStudy (amistre64):

r(t) = ( x(t) , y(t) ) r(t) = (1 - t^2, 2t ) r(t) = ( 0 , 2 ) ; when t = ?

OpenStudy (anonymous):

x= 1 - t^2 y= 2t x = 0, y=2 0 = 1 - t^2 2 = 2t t = 1

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

very good :) now we can define our tangent vector with t=1

OpenStudy (anonymous):

T=<-2,2> i dont know what to do now :S

OpenStudy (anonymous):

i know the point that is (0,2) the tangent vector is <-2,2>

OpenStudy (amistre64):

oh but you do :) scale that vector and add it to our point like you did before. But this time we can sepearte it out so it looks more conventional

OpenStudy (amistre64):

(0,2) + s<-2,2> is cute and all, but not quite standard format

OpenStudy (aravindg):

nikvIst can u help?

OpenStudy (anonymous):

x = 0 -s2 y = 2 + s2?

OpenStudy (amistre64):

yes, but lets clean it up some :) x = -2s y = 2+2s this is parametric, we could define it like normal since we started out in 2 planes

OpenStudy (amistre64):

x = -2s; s = -x/2 y = 2 + 2s; y = 2 + 2(-x/2) y = 2-x

OpenStudy (anonymous):

:O owhh i see... thats the tangent line!

OpenStudy (amistre64):

it is :)

OpenStudy (anonymous):

you're really a hero :)

OpenStudy (amistre64):

thnx ;)

OpenStudy (anonymous):

oke.. one question could you help me with the gradient derivative?

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

i almost have the answer of it, but the last step

OpenStudy (anonymous):

gradient f(2,2) = 7i-2j

OpenStudy (anonymous):

now i have to calculate the direction of the gradient

OpenStudy (amistre64):

lets start with the problem from the beginning so I can see where we are going ... or is this the beginnning?

OpenStudy (anonymous):

\[\sqrt{7^{2} + 2^{2}} = \sqrt{53}\]

OpenStudy (amistre64):

direction has to deal with trig, it is the tan-1(y,x)

OpenStudy (anonymous):

owhh oke.. i have to find the maximum direction at point (2,2)

OpenStudy (amistre64):

maximum direction; would that mean magnitude as well as direction?

OpenStudy (anonymous):

i already calculated the gradient of f and that is (7i-2j)

OpenStudy (amistre64):

the gradient is the normal vector and we can regard if as simply: <7,-2>

OpenStudy (anonymous):

correct :)

OpenStudy (amistre64):